You are designing a rectangular birthday card for a friend. You want the card's length to be 1 in. more than twice the card's width. The area of the card is 88 in.^2. What are the dimensions of the card? Round to the nearest tenth if necessary.
Any help is appreciated.

Answer by stanbon(75887)

You are designing a rectangular birthday card for a friend. You want the card's length to be 1 in. more than twice the card's width. The area of the card is 88 in.^2. What are the dimensions of the card? Round to the nearest tenth if necessary.
----
Width = x
length 2x+1
---
Equation:
x(2x+1) = 88
2x^2 + x - 88 = 0
----
x = [-1 +- sqrt(1-4*2*-88)}/4
----
x = [-1 +- sqrt(1+704)]/4
----
x = [-1 +- 26.55]/4
----
Positive solution:
x = 25.55/5
---
x = 5.11 inches
x+1 = 6.11 inches
====================
Cheers,
Stan H.
======================